OpenStudy (anonymous):

evaluate the expression for r = 2 and t = 7 3(r^2+ t) - r ^2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

3*(4+7)-4=33-4=?????????

OpenStudy (anonymous):

30 lol

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

oops its subtraction lol i meant 29

OpenStudy (alexandervonhumboldt2):

29

OpenStudy (alexandervonhumboldt2):

correct

OpenStudy (anonymous):

i tagged you in a new one